Find an equation of the circle that has center (-1, 1)and passes through(-5,6)
Answer by DrBeeee(684)

Equation of circle is
(1) (x-a)^2 + (y-b)^2 = r^2, where (a,b) are the coordinates of the center of the circle and r is the radius. In your case you have
(2) (x+1)^2 + (y-1)^2 = r^2
To find r substitute the point (-5,6) into (2) and get
(3) (-5+1)^2 + (6-1)^2 = r^2 or
(4) 16 + 25 = r^2 or
(5) r^2 = 41
The equation of the circle is
(6) (x+1)^2 + (y-1)^2 = 41